OpenStudy (anonymous):

* Need help please! How would you graph the following?* 1/2 x+ 7>3 thank u!!

OpenStudy (anonymous):

i bet there is a typo in the question

OpenStudy (anonymous):

\(\frac{1}{7}+7=7\tfrac{1}{2}\) is just a number, and it is not larger than 3

OpenStudy (anonymous):

\[\frac{ 1 }{ 2 }+7=7\frac{ 1 }{2} is~ a ~number~ and~ larger~ than~ 3\]

OpenStudy (anonymous):

I fixed the question. so sorry!

OpenStudy (anonymous):

i had left out the variable x

OpenStudy (superdavesuper):

 it should look something like this :)

OpenStudy (anonymous):

hmmm